i forgot how to do slope intercept form because point slope confused me this week. My question is how do you write an equation in slope intercept with an equation like (3,1); slope=2
Answer by nerdybill(7384)

Plug the given info:
(3,1); slope=2
.
into the "point-slope" form of
y - y1 = m(x - x1)
y - 1 = 2(x - 3)
y - 1 = 2x - 6
y = 2x - 5 (this is what they're looking for)